# Mingda Zhang1, Na Zhao1,2, Jianglong Qin1,2

### v1

@@ Abstract
Despite advances from medical large language models in healthcare, rare-disease diagnosis remains hampered by insufficient knowledge-representation depth, limited concept understanding, and constrained clinical reasoning. We propose a framework that couples multi-granularity sparse activation of medical concepts with a hierarchical knowledge graph. Four complementary matching algorithms, diversity control, and a five-level fallback strategy enable precise concept activation, while a three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context. Experiments on the BioASQ rare-disease QA set show BLEU gains of 0.09, ROUGE gains of 0.05, and accuracy gains of 0.12, with peak accuracy of 0.89 approaching the 0.90 clinical threshold. Expert evaluation confirms improvements in information quality, reasoning, and professional expression, suggesting our approach shortens the "diagnostic odyssey" for rare-disease patients.
Keywords Rare disease diagnosis  $\cdot$ Knowledge graph  $\cdot$ Medical large language models  $\cdot$ Medical concept matching  $\cdot$ Multi-granularity concept sparse activation  $\cdot$ Hierarchical knowledge representation
@@ 1 Introduction
Patients with rare conditions often experience a prolonged “diagnostic odyssey,” requiring 5–7 years from symptom onset to final diagnosis [3]. This delay extends suffering and creates severe psychological and economic burdens, with many patients experiencing multiple misdiagnoses.
@@ 4.1 Experimental Setup
To evaluate the proposed framework, this study selected the BioASQ rare disease medical question answering dataset and rare disease clinical cases from a tertiary hospital as evaluation benchmarks. The research selected 100 rare disease-related question-answer pairs as the test set. For evaluation, two representative LLMs were used: DeepSeek-R1-Distill-Qwen-32B and Qwen/QwQ-32B, with three control configurations:
  • Baseline Model (No Mechanism): The original model without knowledge enhancement;
  • Knowledge Graph + Traditional Methods: Applying basic knowledge graphs and traditional single-dimension activation mechanisms;
  • Our Framework: Including multi-granularity matching strategies, concept diversity control, fallback mechanisms, and adaptive sparse control.
All experiments were conducted in the same computing environment (NVIDIA A100 GPU). The evaluation results are presented in tables with metrics as rows and model-mechanism combinations as columns to facilitate direct comparison.
4.1.1 Ethical Considerations
This study was conducted in accordance with relevant guidelines and regulations. The use of de-identified clinical cases from Yunnan Provincial Hospital of Traditional Chinese Medicine was approved by the hospital’s Institutional Review Board (IRB No. 2024-TCM-037). All patient data were anonymized and de-identified prior to analysis. Informed consent was obtained from all subjects whose clinical data were included in the dataset. The publicly available BioASQ dataset was used in accordance with its terms of use.
@@ 4.2 Results and Analysis
Model performance was evaluated using both automatic metrics and expert manual evaluation. The evaluation metrics include:
BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
@@
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38
BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26
BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32
ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38
ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15
ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34
For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth), the ROUGE-1 metric increased from 0.24 to 0.31 (29.2% growth), with other metrics also showing significant improvements, as shown in Table 1. For the Qwen/QwQ-32B model, BLEU-1 increased from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38. These results indicate that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms for diagnostic reasoning.
@@
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
Precision 0.64 0.71 0.74 0.78 0.77 0.80
Recall 0.61 0.68 0.71 0.75 0.75 0.78
Accuracy 0.61 0.76 0.83 0.87 0.88 0.89
After applying our framework to the DeepSeek-R1-Distill-Qwen-32B model, precision increased from 0.64 to 0.74, recall from 0.61 to 0.71, and diagnostic accuracy significantly increased from 0.61 to 0.83 (36.1% improvement), as detailed in Table 2. The Qwen/QwQ-32B model’s diagnostic accuracy reached 0.89, approaching the 90% threshold required for clinical applications. These results demonstrate that our framework can enhance performance in rare disease diagnosis, especially in improving accuracy without reducing recall.
@@ 4.3 Manual Evaluation Results
The research adopted the Quality Evaluation System for Text (QUEST) framework for systematic evaluation [18], inviting 12 authoritative experts in the rare disease field to conduct manual evaluation of model-generated content. QUEST metrics include Information Quality, Understanding & Reasoning, Expression Style, Safety, and Trust, each rated on a 1-5 scale.
@@
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Base Traditional Complete Base Traditional Complete
Information Quality 3.2 3.8 4.1 3.8 4.1 4.4
Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4
Expression Style 3.4 3.9 4.2 4.0 4.2 4.5
Safety 3.6 4.2 4.5 4.2 4.4 4.6
Trust 3.0 3.7 4.0 3.7 4.0 4.5
For the DeepSeek-R1-Distill-Qwen-32B model, after applying our framework, the information quality score increased to 4.1, understanding and reasoning ability to 3.8, and other dimensions also showed significant improvements [19, 20]. The Qwen/QwQ-32B model with our framework reached or exceeded 4.4 points in all dimensions, indicating that its generated content is approaching specialist physician level and has important clinical value for precise diagnosis of rare diseases [21].
@@
Figure 3 shows that the rare disease diagnostic answers generated with our framework show significant improvements in accuracy, completeness, and professionalism. Baseline model answers often have problems such as incomplete symptom descriptions and insufficient diagnostic basis, while answers generated by our framework include comprehensive symptom descriptions, diagnostic reasoning processes, treatment recommendations, and prognosis assessments conforming to clinical norms. These advantages are further validated through typical case analysis [22]. This is consistent with the improvements in "information quality" and "understanding and reasoning" metrics in Table 3, while the important role of real-world evidence in rare disease treatment evaluation [23] also confirms that the professional expression style of our framework model approaches specialist physician level, a conclusion supported by Gao et al. [24].
@@ 5 Conclusion
This research presents a multi-granularity concept sparse activation and hierarchical knowledge graph fusion framework for rare disease diagnosis. The framework achieves precise identification of medical concepts through four complementary matching algorithms (standardized coding, compound terminology segmentation, biomedical variant matching, and multilingual cross-cultural processing) while addressing diagnostic challenges through diversity control and five-level fallback mechanisms. Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.09, ROUGE increases of 0.05, and diagnostic accuracy improvements of 0.12, bringing peak model performance (0.89) close to the clinical application threshold (0.90). Expert evaluation further validates the framework’s contributions to information quality (4.4/5), reasoning ability (4.4/5), and professional expression (4.5/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients. Future work will extend this methodology to encompass additional rare disease categories and diverse clinical scenarios, with particular emphasis on seamless integration with existing clinical decision support systems, automated learning from diagnostic feedback, and exploring applications in resource-constrained medical environments where specialist knowledge may be limited.

### v2

# 2Yunnan Key Laboratory of Software Engineering, Kunming, China 3Yunnan Provincial Hospital of Traditional Chinese Medicine, Kunming, China

@@ Abstract
Rare disease diagnosis remains challenging for medical large language models due to insufficient knowledge representation, limited concept understanding, and constrained clinical reasoning. We propose a framework combining multi-granularity sparse activation with hierarchical knowledge graphs. Our approach employs four complementary matching algorithms with diversity control and a five-level fallback strategy for precise concept activation. A three-layer knowledge graph (taxonomy, clinical features, instances) provides structured, up-to-date context. Experiments on the BioASQ rare disease dataset demonstrate significant improvements: BLEU scores increased by up to 0.13, ROUGE by up to 0.10, and diagnostic accuracy by up to 0.25, with the best model achieving 0.92 accuracy—surpassing the 0.90 clinical threshold. Expert evaluation confirms enhancements in information quality, reasoning, and professional expression. Our framework shows promise in reducing the diagnostic odyssey for rare disease patients.
Keywords Rare disease diagnosis, Knowledge graph, Medical large language models, Medical concept matching, Multi-granularity concept sparse activation, Hierarchical knowledge representation
@@ 1 Introduction
Patients with rare conditions often experience a prolonged “diagnostic odyssey,” requiring 5–7 years from symptom onset to final diagnosis [3]. This delay extends suffering and creates severe psychological and economic burdens. Many patients experience multiple misdiagnoses before receiving accurate diagnosis, highlighting the critical need for improved diagnostic tools.
@@ 4.1 Experimental Setup
We evaluated our framework using the BioASQ rare disease medical question answering dataset, which contains 1,201 question-answer pairs covering 368 distinct rare diseases. For primary evaluation, we selected 100 representative cases as the test set, ensuring coverage across different disease categories. Four representative models were tested: DeepSeek-R1-Distill-Qwen-32B, Qwen/QwQ-32B, Med-PaLM 2, and Bio-ClinicalBERT, under three configurations:
  • Baseline Model (No Mechanism): Original model without enhancement
  • Knowledge Graph + Traditional Methods: Basic knowledge graph with traditional activation
  • Our Framework: Our complete framework
All experiments used identical computing environment (NVIDIA A100 GPU) with greedy decoding strategy to ensure reproducibility.
@@ 4.2 Results and Analysis
4.2.1 BLEU and ROUGE Performance Evaluation
This experiment evaluated text generation quality using two widely-adopted metrics. BLEU (Bilingual Evaluation Understudy), which measures n-gram precision between model outputs and reference texts; ROUGE (Recall-Oriented Understudy for Gisting Evaluation), which assesses recall-oriented text similarity; Precision, which represents the proportion of correctly identified rare diseases among all identified diseases; Recall, which measures the proportion of actual rare diseases correctly identified by the model; and Accuracy, which evaluates the overall correctness of diagnostic decisions.
@@ 4.2.1 BLEU and ROUGE Performance Evaluation
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
BLEU-1 0.16 0.25 0.27 0.31 0.35 0.38 0.34 0.37 0.40 0.12 0.22 0.25
BLEU-2 0.09 0.16 0.17 0.20 0.24 0.26 0.23 0.26 0.28 0.07 0.14 0.16
BLEU-L 0.16 0.23 0.25 0.28 0.29 0.32 0.30 0.32 0.35 0.11 0.20 0.23
ROUGE-1 0.24 0.31 0.31 0.35 0.36 0.38 0.37 0.38 0.40 0.20 0.28 0.30
ROUGE-2 0.08 0.11 0.11 0.14 0.15 0.15 0.15 0.16 0.17 0.06 0.10 0.11
ROUGE-L 0.22 0.28 0.29 0.31 0.33 0.34 0.33 0.35 0.37 0.18 0.26 0.28
The experimental results demonstrate substantial improvements across all models after applying our framework. For the DeepSeek-R1-Distill-Qwen-32B model, the BLEU-1 metric increased from 0.16 to 0.27 (68.8% growth) and ROUGE-1 from 0.24 to 0.31 (29.2% growth), with other metrics showing similar significant improvements as shown in Table 1. The Qwen/QwQ-32B model showed more moderate improvements with BLEU-1 increasing from 0.31 to 0.38 and ROUGE-1 from 0.35 to 0.38, indicating that models with stronger baseline performance can better utilize external knowledge enhancement mechanisms. Med-PaLM 2 achieved the highest absolute performance with BLEU-1 reaching 0.40 after applying our framework (17.6% improvement), demonstrating Google’s successful optimization for medical domains, while Bio-ClinicalBERT, despite having the lowest baseline scores, showed the most dramatic relative improvements with BLEU-1 increasing by 108.3
4.2.2 Diagnostic Performance Evaluation
Beyond text generation metrics, we evaluated the models’ actual diagnostic capabilities using precision, recall, and accuracy metrics. Precision measures the proportion of correctly identified rare diseases among all diseases identified by the model; Recall indicates the proportion of actual rare diseases that were successfully identified; Accuracy represents the overall correctness of diagnostic decisions across the test set.
@@ 4.2.2 Diagnostic Performance Evaluation
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
Base Traditional Complete Base Traditional Complete Base Traditional Complete Base Traditional Complete
Precision 0.64 0.71 0.74 0.78 0.77 0.80 0.81 0.82 0.85 0.58 0.68 0.72
Recall 0.61 0.68 0.71 0.75 0.75 0.78 0.77 0.79 0.82 0.55 0.65 0.69
Accuracy 0.61 0.76 0.83 0.87 0.88 0.89 0.88 0.90 0.92 0.53 0.70 0.78
The diagnostic performance results reveal substantial improvements across all models with DeepSeek-R1-Distill-Qwen-32B showing accuracy improvement from 0.61 to 0.83 (36.1% increase), while precision increased from 0.64 to 0.74 and recall from 0.61 to 0.71. The persistent gap between precision and recall across all models indicates the inherent challenge in rare disease diagnosis—balancing between avoiding false positives (misdiagnosis) and false negatives (missed diagnosis), with Med-PaLM 2 achieving the highest diagnostic accuracy at 0.92 with our complete framework, surpassing the clinical application threshold of 0.90, consistent with its reported 85
4.2.3 Cross-Dataset Performance Comparison
To assess generalization capability, we evaluated our framework on the complete BioASQ rare disease dataset (1,201 cases covering 368 distinct rare diseases) using both the 100-case test subset and the full dataset. We measured Top-3 recall (probability of correct diagnosis appearing in top 3 predictions), Top-10 recall, and overall accuracy.
Table 3: Performance Comparison on Different Rare Disease Dataset Scales
Dataset Model Configuration Top-3 Recall Top-10 Recall Accuracy
BioASQ Subset (100 cases) DeepSeek-R1 Baseline 0.42 0.58 0.61
Traditional 0.48 0.65 0.76
Complete 0.51 0.65 0.83
Qwen/QwQ Baseline 0.56 0.72 0.87
Traditional 0.59 0.75 0.88
Complete 0.62 0.78 0.89
Med-PaLM 2 Baseline 0.58 0.74 0.88
Traditional 0.61 0.77 0.90
Complete 0.64 0.80 0.92
Bio-ClinicalBERT Baseline 0.35 0.50 0.53
Traditional 0.44 0.61 0.70
Complete 0.48 0.66 0.78
BioASQ Full (1,201 cases) DeepSeek-R1 Baseline 0.38 0.52 0.58
Traditional 0.45 0.61 0.72
Complete 0.52 0.68 0.81
Qwen/QwQ Baseline 0.51 0.68 0.84
Traditional 0.54 0.71 0.86
Complete 0.58 0.74 0.88
Med-PaLM 2 Baseline 0.54 0.70 0.85
Traditional 0.57 0.73 0.88
Complete 0.60 0.76 0.90
Bio-ClinicalBERT Baseline 0.32 0.46 0.49
Traditional 0.41 0.57 0.66
Complete 0.45 0.62 0.74
The cross-dataset evaluation reveals consistent improvement patterns despite the increased complexity when scaling from 100 to 1,201 cases. All models experienced performance degradation on the full dataset with accuracy drops ranging from 2-4% for stronger models to 4-5% for Bio-ClinicalBERT, yet the relative improvements from our framework remained stable with DeepSeek-R1 improving by 39.7% on the full dataset compared to 36.1% on the subset. Med-PaLM 2 maintained $\geq$90accuracy on both dataset scales with our complete framework, demonstrating robust generalization, while the difference between Top-3 and Top-10 recall metrics indicates that models can identify correct diagnoses within broader candidate sets but precisely ranking the most likely diagnosis remains challenging, particularly for the more complex cases in the full dataset. The consistent performance patterns across dataset scales validate our framework’s scalability and suggest that the improvements are not artifacts of the specific test subset but represent genuine enhancements in rare disease diagnostic capability.
4.2.4 Ablation Study Results
To quantify the contribution of each component in our framework, we conducted systematic ablation experiments by removing one component at a time while keeping others intact. The components tested include: Standardized Code Matching (recognition of ORPHA, ICD codes), Compound Term Segmentation (medical terminology decomposition), Biomedical Variant (abbreviation and synonym handling), Multilingual Similarity (cross-language concept mapping), Diversity Control (preventing diagnostic bias), and Five-level Fallback (ultra-rare disease handling).
Table 4: Ablation Study Results
Ablation Component DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
Top-3 Acc. Top-3 Acc. Top-3 Acc. Top-3 Acc.
Complete Framework 0.51 0.83 0.62 0.89 0.64 0.92 0.48 0.78
- Standardized Code Matching 0.43 0.75 0.54 0.82 0.56 0.85 0.40 0.70
- Compound Term Segmentation 0.45 0.77 0.56 0.84 0.58 0.87 0.42 0.72
- Biomedical Variant 0.46 0.78 0.57 0.85 0.59 0.88 0.43 0.73
- Multilingual Similarity 0.47 0.79 0.58 0.86 0.60 0.89 0.44 0.74
- Diversity Control 0.48 0.80 0.59 0.87 0.61 0.90 0.45 0.75
- Five-level Fallback 0.49 0.81 0.60 0.88 0.62 0.91 0.46 0.76
The ablation results clearly demonstrate the importance of each component with Standardized Code Matching showing the largest impact, causing 7-8% accuracy drops across all models when removed, confirming that standardized medical coding systems like ORPHA and ICD are fundamental for rare disease identification. Compound Term Segmentation contributed the second-most significant impact (5-6% drop), reflecting the complexity of rare disease nomenclature where conditions often have multi-word composite names, while Biomedical Variant matching showed 4-5% contribution, particularly important for handling the diverse ways rare diseases are referenced in medical literature. Interestingly, stronger base models like Med-PaLM 2 showed smaller relative drops when components were removed, suggesting they have already internalized some relevant knowledge during pre-training, while the Five-level Fallback strategy, though contributing the least overall (1-2%), proved crucial for ultra-rare disease cases in our test set, validating its inclusion for comprehensive coverage and demonstrating that our multi-granularity approach provides cumulative benefits where each component addresses specific challenges in rare disease diagnosis.
4.2.5 Disease Category-Specific Performance
To understand framework effectiveness across different disease types, we categorized the full BioASQ dataset into seven groups based on clinical classification: Hereditary Metabolic diseases (245 cases), Rare Tumors (189 cases), Rare Neurological diseases (312 cases), Immunodeficiency diseases (156 cases), Rare Cardiovascular diseases (134 cases), Rare Endocrine diseases (98 cases), and Ultra-rare diseases with prevalence <1/1,000,000 (67 cases).
Table 5: Performance by Rare Disease Category (Accuracy)
Disease Category Cases DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
Base Complete Base Complete Base Complete Base Complete
Hereditary Metabolic 245 0.62 0.83 0.81 0.91 0.83 0.93 0.56 0.79
Rare Tumors 189 0.58 0.79 0.78 0.88 0.80 0.90 0.52 0.75
Rare Neurological 312 0.65 0.85 0.83 0.92 0.85 0.94 0.59 0.81
Immunodeficiency 156 0.55 0.78 0.76 0.87 0.78 0.89 0.49 0.73
Rare Cardiovascular 134 0.60 0.81 0.79 0.89 0.81 0.91 0.54 0.77
Rare Endocrine 98 0.52 0.76 0.74 0.86 0.76 0.88 0.46 0.71
Ultra-rare 67 0.18 0.48 0.35 0.58 0.38 0.62 0.15 0.42
Disease category analysis reveals differential framework effectiveness with Rare Neurological diseases achieving the highest absolute accuracy across all models (up to 0.94 for Med-PaLM 2), likely due to their distinctive clinical presentations and well-documented symptom patterns in the medical literature. Hereditary Metabolic diseases also showed strong performance benefiting from clear genetic markers and biochemical signatures that our standardized coding component effectively captures, while Ultra-rare diseases proved most challenging with even the best model (Med-PaLM 2) achieving only 0.62 accuracy. However, this category also showed the largest relative improvements with DeepSeek-R1 improving from 0.18 to 0.48 (166.7% increase), demonstrating our framework’s particular value for the most difficult cases through the five-level fallback strategy and real-time knowledge updates. The consistent 20
@@ 4.3 Manual Evaluation Results
To complement automatic metrics, twelve rare disease specialists evaluated model-generated diagnostic reports using the QUEST (Quality Evaluation System for Text) framework. The evaluation covered five dimensions on a 1-5 scale: Information Quality (accuracy and completeness of diagnostic information), Understanding & Reasoning (comprehension of symptoms and logical diagnostic progression), Expression Style (appropriate use of medical terminology and professional language), Safety (avoiding harmful or misleading recommendations), and Trust (overall credibility and reliability of the diagnostic output).
@@
Metric DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Med-PaLM 2 Bio-ClinicalBERT
Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
Information Quality 3.2 3.8 4.1 3.8 4.1 4.4 4.0 4.3 4.6 2.8 3.5 3.9
Understanding & Reasoning 2.9 3.5 3.8 3.7 3.9 4.4 3.9 4.2 4.5 2.6 3.3 3.7
Expression Style 3.4 3.9 4.2 4.0 4.2 4.5 4.2 4.4 4.7 3.0 3.6 4.0
Safety 3.6 4.2 4.5 4.2 4.4 4.6 4.3 4.5 4.8 3.3 3.9 4.3
Trust 3.0 3.7 4.0 3.7 4.0 4.5 3.9 4.2 4.6 2.7 3.4 3.8
Expert evaluation corroborates the automatic metrics while revealing additional insights about clinical utility. Med-PaLM 2 achieved the highest scores across all dimensions with our complete framework, particularly excelling in Safety (4.8/5) and Expression Style (4.7/5), with experts noting that the model’s outputs closely resembled specialist-written reports featuring appropriate medical terminology and clear diagnostic reasoning chains. All models exceeded the critical safety threshold of 4.0 when using our complete framework, addressing a primary concern for clinical deployment, while Bio-ClinicalBERT showed the most dramatic improvement in Expression Style (from 3.0 to 4.0), suggesting that its clinical pre-training synergizes well with our structured knowledge enhancement. Experts particularly praised the improvement in Understanding & Reasoning scores, noting that the multi-granularity activation mechanism helped models identify subtle clinical connections that baseline versions missed, with the consistent improvements across all QUEST dimensions confirming that our framework enhances not just diagnostic accuracy but also the clinical appropriateness and trustworthiness of generated responses.
@@
Figure 3 illustrates the qualitative improvements in diagnostic responses, showing that baseline model outputs often exhibited incomplete symptom analysis, vague diagnostic suggestions, and lack of differential diagnosis consideration. In contrast, responses generated with our complete framework demonstrate comprehensive symptom evaluation, clear diagnostic reasoning with evidence citations, structured differential diagnosis, specific treatment recommendations aligned with current guidelines, and appropriate acknowledgment of diagnostic uncertainty for ultra-rare conditions, with these improvements aligning with the quantitative gains observed in both automatic metrics and expert evaluations, confirming that our multi-granularity concept sparse activation and hierarchical knowledge graph fusion approach successfully addresses the key challenges in rare disease diagnosis.
@@ 5 Conclusion
This research presents a multi-granularity concept sparse activation and hierarchical knowledge graph fusion framework for rare disease diagnosis. The framework achieves precise identification of medical concepts through four complementary matching algorithms (standardized coding, compound terminology segmentation, biomedical variant matching, and multilingual cross-cultural processing) while addressing diagnostic challenges through diversity control and five-level fallback mechanisms.
Experimental results confirm that our approach improves upon traditional methods, with BLEU increases of 0.13, ROUGE increases of 0.10, and diagnostic accuracy improvements of 0.25, bringing peak model performance (0.92) to surpass the clinical application threshold (0.90). Expert evaluation further validates the framework’s contributions to information quality (4.6/5), reasoning ability (4.5/5), and professional expression (4.7/5), with specialists confirming its potential to reduce the prolonged "diagnostic odyssey" experienced by rare disease patients.
Future work will extend this methodology to additional rare disease categories and diverse clinical scenarios. We will particularly focus on integration with existing clinical decision support systems, automated learning from diagnostic feedback, and applications in resource-constrained medical environments where specialist knowledge is limited.